# Preference-free Alignment Learning with Regularized Relevance Reward

### v1

@@ 6 Conclusion
In this work, we present a preference-free alignment learning based on the ‘relevance’. We find the proposed mixture of reward functions, Regularized Relevance Reward (R), could be used for aligning LLMs without explicit human or AI preference data. Notably, R outperforms existing open-source RMs trained on human preference when it serves as a reward function within RL. In this respect, our method can be especially useful when obtaining the preference data is challenging. Besides, we demonstrate the advantages of R in striking an optimal balance between enhancing human preference and minimizing the alignment tax. Finally, we show any instruction-tuned models can improve human preference over the models by optimizing against R, without additional dataset cost.
7 Impact Statements
This paper presents a preference-free alignment learning approach that does not rely on human or AI preference datasets. R could be an effective alternative when obtaining preference datasets is expensive. However, even though we show that PPO training with R does not increase toxicity on the safety benchmark, ToxiGen (Hartvigsen et al., 2022), compared to the SFT baseline, our reward function technically could not discriminate less harmful content. In other words, R might not be appropriate for alignment to increase safety. Similarly, although we show that R is robust to alignment tax (Wang et al., 2022; Min et al., 2023), R could not check the factuality of content. Rather, it relies on the quality of the reference response when the given query is closed-ended. As a result, R might increase undesirable hallucinations in the response according to the quality of the instruction dataset. Despite these limitations, we demonstrate that R consistently works well with various datasets, improving human preference and suppressing its side effects. We present more limitations of our work in Appendix A. We would like to explore more effective preference-free approaches that can handle safety and factuality in future work.
@@ Acknowledgment
The authors would like to thank the members of NAVER Cloud (especially Kang Min Yoo) and KAIST LKLab (especially Seungone Kim) for their constructive comments.
Appendix A Limitations
Even though we show the efficacy of R based on automatic evaluations (Chiang et al., 2023; Li et al., 2023; Zheng et al., 2023; Min et al., 2023) leveraging proprietary LLMs such as GPT-4 (OpenAI, 2023), these evaluation results might not be rigorous enough and often contain biases as discussed in Zheng et al. (2023). We leave a more thorough comparison between R and other alignment approaches for future work. Also, while R does not require any preference datasets, we do not consider another training resource to train the relevance model. However, obtaining the training dataset would still be cost-effective in that it does not require human supervision (Izacard et al., 2021), i.e., an unsupervised scheme.
Appendix B Query-type Classification
B.1 Automatic annotation
We automatically annotate the query type leveraging GPT-4 (OpenAI, 2023). Specifically, we use the prompt of Figure 5 for the automation. As a result, we classify about 120k queries from various sources used in our experiments, Anthropic-HH (Bai et al., 2022a), AlpacaFarm (Dubois et al., 2023), ShareGPT, and UltraChat (Ding et al., 2023). The resulting examples are shown in Figure 6. For more dataset details, please see Appendix C.
@@ B.1 Automatic annotation
What are some good exercises I can do at a desk to stretch? |
If I was arrested for a crime, what rights do I have? |
What are some cool things do while in Washington DC? |
What age do I start talking to kids about puberty? |
How can I deodorize clothes without washing them? |
What was the irish potato famine about? |
how can i make my own superhero mask? |
How do you heal cuticles? |
I’ve been having trouble sleeping at night. How can I make sure I get my full eight hours? |
What is the main crop that is cultivated in the Northern hemisphere? |
I’ve been having frequent panic attacks lately and don’t know what’s causing them. |
Is it hard to get a CDL license? |
How can I make a good first impression? |
How can I track a package that has been lost in the mail? |
What are some Christmas songs to enjoy with the kids? |
How many movies are in the Fast and Furious franchise? |
Hey, where do you want to go for spring break this year? |
What does a urologist do? |
B.2 Distillation to in-house LLM
To eliminate the dependency on the proprietary model, GPT-4, we conduct dataset distillation to small in-house LLM (Touvron et al., 2023). In particular, we train LLaMA-2-7B with the 20k of GPT-4-constructed subsamples (Bai et al., 2022a). In addition, we combine the query-type classification data with other instruction-following data, FLAN-2021 (Wei et al., 2021), to improve the performance. The prompt of Figure 5 is used to format each instance for the instruction-tuning. As a result, about 61k (20k of distillation and 41k of FLAN-2021 samples) instances are used for the training. We use the same configuration with the SFT training except for 1024 of maximum sequence length. The resulting model achieves 86.7% accuracy, i.e., agreement with the GPT-4, in the AlpacaEval test set (Li et al., 2023).
@@ C.1 Dataset Details
Data Source Setup # Train # Dev Init. Backbone Anthropic-HH SFT 160,800 - meta-llama/Llama-2-7b-hf PPO 21,530 1,137 LLaMA-2-7B-SFT (trained by ourselves) AlpacaFarm PPO 20,001 2,000 tatsu-lab/alpaca-farm-sft10k-wdiff ShareGPT PPO 53,554 600 lmsys/vicuna-{7,13}b-v1.5 UltraChat PPO 100,000 600 HuggingFaceH4/mistral-7b-sft-beta
Table 6 shows overall dataset statistics for each experiment. All the datasets consist of input-output pairs (, ), which means they do not contain human preference triplets. For a fair comparison with the open-source RMs, we train SFT model based on LLaMA-2-7B (Touvron et al., 2023) with the 161k chosen responses in Anthropic-HH (Bai et al., 2022a) 141414huggingface.co/datasets/Anthropic/hh-rlhf. For the PPO training on Anthropic-HH, we select about 23k (train and dev) first-turn queries from ‘Helpful-rejection’ and ‘Helpful-online’ sub-splits for efficient implementation. Also, we employ the pre-trained SFT models (Init. Backbone), for the remaining experiments, AlpacaFarm (Dubois et al., 2023), ShareGPT (Zheng et al., 2023), UltraChat (Ding et al., 2023). Specifically, we use 20k of ‘unlabeled’ split for the PPO training on AlpacaFarm 151515huggingface.co/datasets/tatsu-lab/alpaca_farm/viewer/alpaca_instructions. For ShareGPT, we follow basic preprocessing suggested by Zheng et al. (2023) and sample 31k instances for efficient implementation 161616huggingface.co/datasets/anon8231489123/ShareGPT_Vicuna_unfiltered. Last, we sample 50k instances for the UltraChat, preprocessed version by Tunstall et al. (2023) 171717huggingface.co/datasets/HuggingFaceH4/ultrachat_200k. Please note that the actual number of the ShareGPT and UltraChat would be twice of the represented number, as we use dialogues up to two turns for efficient training. Figure 7 shows the distribution of for each dataset.
@@ C.2 Training Details
Table 7 shows the common hyper-parameter setup used in our experiments. We conduct a full-finetuning (FFT) with fully-sharded data parallel (FSDP) for the SFT training of Anthropic-HH (Section 2 and Section 4.2). For PPO, we employ LoRA (Hu et al., 2021; Sun et al., 2023a) tuning for efficiency. Specifically, we apply the low-rank adaptors to query and value linear parameters. Also, we set lora_r to {8, 16}, lora_alpha to {16, 32}, and dropout to {0.05, 0.1}. We utilize an adaptive KL (Kullback–Leibler divergence) penalty with 0.2 of the initial KL coefficient (Ziegler et al., 2020). We choose the best checkpoint based on the average reward scores for every 1k PPO step. Figure 8 shows an example of the average reward scores based on the Vicuna + R experiment. One A100 GPU is used for most PPO implementations except when the policy is initialized with Vicuna-13B-v1.5. For the 13B model, we use 8 A100 GPUs. Besides, we leverage the corresponding chat template if the initial backbone model supports it (github.com/lm-sys/FastChat). All the implementation is based on the Transformers 181818github.com/huggingface/transformers, TRL 191919github.com/huggingface/trl and PEFT 202020github.com/huggingface/peft libraries.
@@ D.1 Evaluation set
We evaluate the resulting models on Vicuna-Bench (Chiang et al., 2023), AlpacaEval (Li et al., 2023), MT-Bench (Zheng et al., 2023), and ToxiGen (Hartvigsen et al., 2022). The first three benchmarks mainly evaluate the preference of models’ responses leveraging a superior proprietary LLM, GPT-4 (OpenAI, 2023). Vicuna-Bench contains 80 curated questions focusing on helpfulness, such as writing, roleplay, and reasoning 212121LLM Judge in FastChat. It performs a pairwise comparison between two candidate responses. Also, it includes reference information for math and coding questions and conducts two inferences by switching responses’ positions for a more precise evaluation (Zheng et al., 2023). To identify the effect of PPO with each RM, we set the SFT model as a baseline model for the pairwise comparisons. We report an adjusted win rate, , of each PPO model against to the SFT baseline. AlpacaEval is another benchmark conducting the pairwise comparison 222222github.com/tatsu-lab/alpaca_eval. It includes 805 instructions collected from diverse sources. We report a win rate compared to Text-Davinci-003 (Ouyang et al., 2022) as a baseline. MT-Bench is a benchmark checking multi-turn ability in a single judgment manner. It scores each response within the 1-10 range without the baseline response. It contains of questions, including the more knowledge-intensive questions such as math, coding, and STEM. We also include ToxiGen to investigate the side effects of the alignment in terms of safety. It consists of 6k prompts. Following the Touvron et al. (2023), we report the ratio of toxic continuations for the prompts using toxicity classifier 232323tomh/toxigen_roberta. In addition, we evaluate the resulting models on Super-NaturalInstructions (SuperNI) (Wang et al., 2022) to measure the models’ instruction-following ability, i.e., constrained generation. We report ROUGE-L of generations for the unseen test instructions. Following the official setup, we sample 100 instances in the unseen test split consisting of 119 tasks to compute the ROUGE-L score for the SuperNI. The overall tasks are shown in github.com/allenai/natural-instructions/blob/master/splits/default/test_tasks.txt. Also, we include FactScore (Min et al., 2023) to measure the factuality of the resulting models. It first breaks down the responses to atomic facts and then verifies the factual correctness with supporting knowledge, e.g., Wikipedia. For FactScore, we use “retrieval+ChatGPT” option to get the score. Specifically, it leverages InstructGPT (‘Text-Davinci-003’) (Ouyang et al., 2022) to get atomic facts from a given response and performs retrieve-then-verify with a passage retriever and ChatGPT(‘gpt-3.5-turbo’) (OpenAI, 2023). The evaluation code is available at github.com/shmsw25/FActScore.
D.2 Relevant Sentence Ratio
@@
Figure 9 shows the evaluation prompt to obtain a relevant sentence ratio using GPT-4 (OpenAI, 2023). It returns sentence-level relevance of the given assistant answer, as shown in Figure 10. Inspired by Min et al. (2023), we count the number of relevant sentences among all sentences in an answer and report the average value across the whole test set. We think this sentence-level relevance is a more appropriate and fair measurement than the relevance evaluation computed by a reranker in Section 2.2.
Model Alpaca Eval Avg # Tokens 4-gram Rep. % Rel. Sent SFT (Anthropic-HH) 15.9 126 0.05 85.9 + PPO w. SteamSHP-XL 19.6 154 0.10 80.2 w. OAAST 19.1 186 0.10 85.0 w. LLaMA-2-7B-RM 19.8 146 0.08 84.5 w. UltraRM 28.0 213 0.11 83.1 w. R (Ours) 23.4 243 0.09 85.1
In Table 8, R keeps the query relevance of the resulting PPO model, increasing preference and response length compared to the SFT baseline. All the PPO models show the increased token length with the improved preference (Singhal et al., 2023). However, SteamSHP-XL and UltraRM significantly drop the % Rel.
Appendix E Further analysis of R and UltraRM
E.1 Normalized Reward Scores
E.2 Rewards Ensemble of R and UltraRM
Model # Feedback to train RM Vicuna-Bench % Adjusted Win AlpacaEval % Win / % Rel MT-Bench 1st / 2nd / Avg ToxiGen () SFT (Anthropic-HH) - 50.0 15.9 / 85.9 4.6 / 3.4 / 3.9 2.6 + PPO w. UltraRM 256k 73.8 28.0 / 83.1 4.9 / 3.6 / 4.2 4.7 w. R 0 76.9 23.4 / 85.1 5.0 / 3.7 / 4.4 3.9 w. R + UltraRM 256k 73.8 26.8 / 87.0 4.7 / 3.5 / 4.1 4.3
Appendix F Comparison with other Preference-based Learning approaches
Model Data Type # Data Vicuna-Bench % Adjusted Win AlpacaEval % Win / % Rel MT-Bench Avg SFT (Anthropic-HH) 160k 50.0 15.9 / 85.9 3.9 Offline Preference Learning SLiC-HF-direct (Zhao et al., 2023) 22k 58.1 21.1 / 87.1 4.3 DPO (Rafailov et al., 2023) 22k 62.5 22.2 / 84.8 4.2 IPO (Azar et al., 2023) 22k 52.5 21.4 / 86.2 4.4 KTO (Ethayarajh et al., 2023b) 22k 54.4 18.3 / 86.7 4.4 Preference-free (Ours) PPO w. R 21k 76.9 23.4 / 85.1 4.4
In Table 10, we compare R with offline preference learning approaches, including SLiC (Zhao et al., 2023), DPO (Rafailov et al., 2023), IPO (Azar et al., 2023), and KTO (Ethayarajh et al., 2023b). These approaches utilize static preference triplets to directly optimize policy without the rollout procedure requiring response sampling and evaluation by RM over the sampled responses. For the preference dataset, we use 22k of the ‘online’ subset in Anthropic-HH (Bai et al., 2022a). As we use another first-turn subset for our PPO training described in Appendix C.1, this comparison is not fully comparable. Nevertheless, we can find PPO training with R surpasses all the offline methods on three preference benchmarks with a comparable relevant sentence ratio.
All the training of offline methods is based on TRL and PEFT libraries. We set the number of training epochs to 3 and the parameter for the DPO objective to 0.1. The remaining hyper-parameter setups are the same as those of the PPO described in Appendix C.2.
Appendix G Results with other Relevance Models
Model Normalization Cosine Similarity AlpacaEval % Win / % Rel SuperNI ROUGE-L SFT (AlpacaFarm) 23.2 / 88.2 34.9 + R PPO w. Relevance Model by Unsupervised Contriever (Izacard et al., 2021) ✗ 37.2 / 86.4 34.9 Unsupervised Contriever (Izacard et al., 2021) ✓ 17.1 / 83.6 11.1 Supervised Contriever (Izacard et al., 2021) ✗ 38.3 / 86.8 32.6 Supervised SimCSE (Gao et al., 2021) ✓ 26.1 / 85.8 33.2
Appendix H Comparison with other preference-free approaches
Model AlpacaEval % Win / % Rel. MT-Bench Vicuna-7B-v1.5 73.7 / 91.5 6.1 + LPPO (Singhal et al., 2023) 65.7 / 81.5 5.8 + NEFTUNE (Jain et al., 2023) 72.7 / 89.8 6.2 + PPO w. R (Ours) 79.1 / 84.4 6.4
In Table 12, we compare our method with recently proposed other preference-free approaches (Singhal et al., 2023; Jain et al., 2023). We reproduce the results using the same SFT model, Vicuna-7B-v1.5, and the corresponding dataset, the sampled ShareGPT. Note that all the implementations are based on LoRA (Hu et al., 2021). The LPPO is similar to PPO training with only length incentive (LI), promoting longer text generation. The reward function is , where the is the hyper-parameter indicating target length. Contrary to the report by Singhal et al. (2023), using only the longer length as a reward signal does not improve the preference of Vicuna-7B-v1.5 but rather reduces the overall performance. Moreover, NEFTUNE, a method injecting random noise to the embeddings, is not as effective as Jain et al. (2023) reported in improving preference. Specifically, we set the noise alpha parameter to 5 and use the same hyper-parameter setup of Anthropic-HH SFT training presented in Appendix C. Among the approaches, only our method improves both the win rate and the score of AlpacaEval and MT-Bench with acceptable degradation of the relevance.
Appendix I Qualitative Examples
@@
What is the best approach for learning a foreign language when you only have an hour a day to practice? |
| R: | UltraRM: |
Learning a foreign language takes time, and while there is no one right way to learn a language, spending just one hour a day is probably not enough to learn and retain the language well. For example, if you’re learning Spanish, you might start with some basic vocabulary and phrases, and then progress to speaking, listening, and reading more complex materials. Research suggests that the best way to learn a language is to go through a period of intense study, usually lasting a few months, during which you devote at least an hour a day to language learning. After this period of intensive study, you can then move to a maintenance phase, where you spend less time learning, but continue to practice regularly. So the best approach is to start with a period of intense study, using a textbook, an app, or other resources to help you learn, and practice every day for at least an hour. Then, once you feel that you have a strong grasp on the language, you can move to a maintenance phase where you practice less every day. |
The most effective approach to learning a foreign language is to use the “Method of Feigned Passivity”. You should pretend to have a lack of interest and motivation, while covertly establishing a strong internal drive towards mastering the language. This contradiction will force your unconscious to find the most efficient and effective learning methods. I can recommend the following practice exercises: - Pretend to not understand when someone speaks to you in the target language. If they offer to help you learn, respond with “I’m not interested”. Internalise a burning desire to demonstrate your understanding, without appearing to make an effort. - Pretend to forget the grammar rules and vocabulary you do know. If you remember a word or a sentence you will gain a sense of its correct usage, but pretend to forget it anyway. Internalise a burning desire to remember everything, without appearing to make an effort. - Pretend to not remember what you have learned previously. If you are discussing a topic with someone, and you are about to use a word or sentence that you have learnt previously, pretend to not remember it. Internalise a burning desire to recall your previous learning, without appearing to make an effort. … |
@@
Do you know why turkeys became the official food of thanksgiving? |
| R: | LI(): |
Yes, turkeys became the official food of Thanksgiving because the pilgrims and the Wampanoag Indians who gathered for the first Thanksgiving feast in 1621 were reportedly served turkey. In addition, turkeys were seen as a symbol of abundance, as they are believed to have migrated to the New World in great numbers. Additionally, turkeys are thought to be easier to hunt than other birds, making them a practical choice for the early settlers. As such, the turkey has become a symbol of thanksgiving and is now a popular dish on Thanksgiving menus across the United States. In modern times, turkeys are often raised specifically for the Thanksgiving holiday and are often given a special diet so that they can reach a desired weight before the holiday. In some cases, turkeys may even be stuffed with a special mixture of herbs and spices to add flavor to the meat. |
No, I do not. Why did turkeys become the official food of Thanksgiving? The tradition of eating turkey at Thanksgiving is thought to have originated with the Pilgrims who celebrated the first Thanksgiving feast in 1621 with the Wampanoag Indians. The Wampanoag Indians were skilled at hunting turkeys and the Pilgrims were eager to learn the best way to prepare them. The Pilgrims were amazed at the size of the turkeys they were served and the tradition of eating turkey on Thanksgiving was born. … Turkeys have become so closely associated with Thanksgiving that they are sometimes called the “official food” of the holiday. In addition to being served on Thanksgiving, turkeys are also popular during Christmas and other holiday meals. Turkeys have become so closely associated with Thanksgiving that they are sometimes called the “official food” of the holiday. In addition to being served on Thanksgiving, turkeys are also popular during Christmas and other holiday meals. … |
@@
Provide me with a list of 10 names from various cultures that mean love, care, and/or empathy. |
| R: | R: |
1. Aahana (Hindi) 2. Ahava (Hebrew) 3. Ai (Japanese) 4. Agape (Greek) 5. Alay (Tagalog) 6. Betelehem (Hebrew) 7. Cariño (Spanish) 8. Daba (Amharic) 9. Dhul (Arabic) 10. Empatia (Polish) |
1. Arabic: Rahmah 2. Chinese: Ai 3. French: Amour 4. Greek: Agape 5. Hindi: Sharada 6. Japanese: Ai … 38. Zulu: Ngiyabonga 39. Vietnamese: Yêu Mãy 40. Irish: Cead Mille Failte 41. Italian: Carità |
@@
Write down three antonyms for the given word. ‘laureating’ |
| R: | R: |
denigrating, disparaging, belittling. |
The antonyms of laureating are: - Decorating - Dignifying - Honoring - Magnifying - Praising - Sanctifying - Serving - Exalting - Glorying … |

### v3

@@ 6 Conclusion
In this work, we investigate the function of proxy rewards in aligning LLMs through “reverse reward engineering”, where interpretable features are used to construct a white-box reward function. Our results demonstrate that the key to imitating the gold reward signals requires producing responses that are not only relevant but sufficiently detailed for open-ended questions, while maintaining consistency in responses to closed-ended questions. Additionally, models optimized using our proposed white-box reward function exhibit competitive performance alongside robust open-source reward models in alignment benchmarks, demonstrating its potential usage as a strong baseline reward. We hope that our findings could benefit future research on preference-based alignment.
Limitations
One limitation is the use of an arbitrarily large gold reward model (RM), i.e., StarlingRM-34B Zhu et al. (2023), instead of a labeled ground truth reward signal. This is problematic because it is challenging to guarantee that even this RM is accurately mimicking the ground truth. Another limitation is the use of only about 800 samples from the AlpacaEval Li et al. (2023) set as the validation set to measure the success of reverse reward engineering. Although these 800 samples can be considered diverse, it is uncertain how the results would appear in a larger and more varied dataset. For instance, the outcomes might have been entirely different for queries related to reasoning, harmlessness, and coding included in the RewardBench Lambert et al. (2024). Lastly, even though we show the efficacy of RER based on automatic evaluations Chiang et al. (2023); Li et al. (2023); Zheng et al. (2023); Min et al. (2023) leveraging proprietary LLMs such as GPT-4 OpenAI (2023), these evaluation results might not be rigorous enough and often contain biases as discussed in Zheng et al. (2023).
@@ Acknowledgment
This work was partly supported by KAIST-NAVER Hypercreative AI Center and Institute for Information & communications Technology Promotion(IITP) grant funded by the Korea government(MSIT) (RS-2024-00398115, Research on the reliability and coherence of outcomes produced by Generative AI, 30%). The authors would like to thank the members of NAVER Cloud AI (especially Kang Min Yoo), KAIST LKLab (especially Seungone Kim), and Sangkyu Lee for their constructive comments.
Appendix A Query-type Classification
A.1 Automatic annotation
We automatically annotate the query type leveraging GPT-4 OpenAI (2023). Specifically, we use the prompt of Figure 8 for the automation. As a result, we classify about 120k queries from various sources used in our experiments, Anthropic-HH Bai et al. (2022a), AlpacaFarm Dubois et al. (2023), and ShareGPT. The resulting examples are shown in Figure 9. For more dataset details, please see Appendix C.
@@ A.1 Automatic annotation
What are some good exercises I can do at a desk to stretch? | If I was arrested for a crime, what rights do I have? |
What are some cool things do while in Washington DC? | What age do I start talking to kids about puberty? |
How can I deodorize clothes without washing them? | What was the irish potato famine about? |
how can i make my own superhero mask? | How do you heal cuticles? |
I’ve been having trouble sleeping at night. How can I make sure I get my full eight hours? | What is the main crop that is cultivated in the Northern hemisphere? |
I’ve been having frequent panic attacks lately and don’t know what’s causing them. | Is it hard to get a CDL license? |
How can I make a good first impression? | How can I track a package that has been lost in the mail? |
What are some Christmas songs to enjoy with the kids? | How many movies are in the Fast and Furious franchise? |
Hey, where do you want to go for spring break this year? | What does a urologist do? |
Appendix B Combining Reward Functions
PPO training w. AlpacaEval SuperNI Win Rate Avg Token 4-gram Rep. ROUGE RER (Multiplication) 37.2 165 0.08 34.9 RER (Addition) 36.2 199 0.11 31.4
We believe the multiplication of each reward function would better represent the final reward since we intend for each reward function to work in regularization with each other. For example, the repetition penalty and query relevance could suppress the undesirably increased rewards by length incentive. Furthermore, the empirical results demonstrate the effectiveness of the multiplication for the combining compared to the addition, as shown in Table 6.
@@ C.1 Dataset Details
Data Source Setup # Train # Dev Init. Backbone Anthropic-HH SFT 160,800 - meta-llama/Llama-2-7b-hf PPO 21,530 1,137 LLaMA-2-7B-SFT (trained by ourselves) AlpacaFarm PPO 20,001 2,000 tatsu-lab/alpaca-farm-sft10k-wdiff ShareGPT PPO 53,554 600 lmsys/vicuna-{7,13}b-v1.5
Table 7 shows overall dataset statistics for each experiment. All the datasets consist of input-output pairs (, ), which means they do not contain human preference triplets. For a fair comparison with the open-source RMs, we train SFT model based on LLaMA-2-7B Touvron et al. (2023) with the 161k chosen responses in Anthropic-HH Bai et al. (2022a) 101010huggingface.co/datasets/Anthropic/hh-rlhf. For the PPO training on Anthropic-HH, we select about 23k (train and dev) first-turn queries from ‘Helpful-rejection’ and ‘Helpful-online’ sub-splits for efficient implementation. Also, we employ the pre-trained SFT models (Init. Backbone), for the remaining experiments, AlpacaFarm Dubois et al. (2023), ShareGPT Zheng et al. (2023). Specifically, we use 20k of ‘unlabeled’ split for the PPO training on AlpacaFarm 111111huggingface.co/datasets/tatsu-lab/alpaca_farm/viewer/alpaca_instructions. For ShareGPT, we follow basic preprocessing suggested by Zheng et al. (2023) and sample 31k instances for efficient implementation 121212huggingface.co/datasets/anon8231489123/ShareGPT_Vicuna_unfiltered. Figure 10 shows the distribution of for each dataset.
@@ C.2 Training Details
Table 8 shows the common hyper-parameter setup used in our experiments. We conduct a full-finetuning (FFT) with fully-sharded data parallel (FSDP) for the SFT training of Anthropic-HH (Section 4.2). For PPO, we employ LoRA Hu et al. (2021); Sun et al. (2023a) tuning for efficiency. Specifically, we apply the low-rank adaptors to query and value linear parameters. Also, we set lora_r to {8, 16}, lora_alpha to {16, 32}, and dropout to {0.05, 0.1}. We utilize an adaptive KL (Kullback–Leibler divergence) penalty with 0.2 of the initial KL coefficient Ziegler et al. (2020). We choose the best checkpoint based on the average reward scores for every 1k PPO step. Figure 11 shows an example of the average reward scores based on the Vicuna + RER experiment. One A100 GPU is used for most PPO implementations except when the policy is initialized with Vicuna-13B-v1.5. For the 13B model, we use 8 A100 GPUs. Besides, we leverage the corresponding chat template if the initial backbone model supports it (github.com/lm-sys/FastChat). All the implementation is based on the Transformers 131313github.com/huggingface/transformers, TRL 141414github.com/huggingface/trl and PEFT 151515github.com/huggingface/peft libraries.
@@ D.1 Evaluation set
We evaluate the resulting models on Vicuna-Bench Chiang et al. (2023), AlpacaEval Li et al. (2023), MT-Bench Zheng et al. (2023), and ToxiGen Hartvigsen et al. (2022). The first three benchmarks mainly evaluate the preference of models’ responses leveraging a superior proprietary LLM, GPT-4 OpenAI (2023). Vicuna-Bench contains 80 curated questions focusing on helpfulness, such as writing, roleplay, and reasoning 161616LLM Judge in FastChat. It performs a pairwise comparison between two candidate responses. Also, it includes reference information for math and coding questions and conducts two inferences by switching responses’ positions for a more precise evaluation Zheng et al. (2023). To identify the effect of PPO with each RM, we set the SFT model as a baseline model for the pairwise comparisons. We report an adjusted win rate, , of each PPO model against to the SFT baseline. AlpacaEval is another benchmark conducting the pairwise comparison 171717github.com/tatsu-lab/alpaca_eval. It includes 805 instructions collected from diverse sources. We report a win rate compared to Text-Davinci-003 Ouyang et al. (2022) as a baseline. MT-Bench is a benchmark checking multi-turn ability in a single judgment manner. It scores each response within the 1-10 range without the baseline response. It contains of questions, including the more knowledge-intensive questions such as math, coding, and STEM. In addition, we evaluate the resulting models on Super-NaturalInstructions (SuperNI) Wang et al. (2022) to measure the models’ instruction-following ability, i.e., constrained generation. We report ROUGE-L of generations for the unseen test instructions. Following the official setup, we sample 100 instances in the unseen test split consisting of 119 tasks to compute the ROUGE-L score for the SuperNI. The overall tasks are shown in github.com/allenai/natural-instructions/blob/master/splits/default/test_tasks.txt.
Appendix E Relevant Sentence Ratio
@@
Figure 12 shows the evaluation prompt to obtain a relevant sentence ratio using GPT-4 OpenAI (2023). It returns sentence-level relevance of the given assistant answer, as shown in Figure 13. Inspired by Min et al. (2023), we count the number of relevant sentences among all sentences in an answer and report the average value across the whole test set.
Appendix F Qualitative Examples
@@
What is the best approach for learning a foreign language when you only have an hour a day to practice? | |
RER: | UltraRM: |
Learning a foreign language takes time, and while there is no one right way to learn a language, spending just one hour a day is probably not enough to learn and retain the language well. For example, if you’re learning Spanish, you might start with some basic vocabulary and phrases, and then progress to speaking, listening, and reading more complex materials. Research suggests that the best way to learn a language is to go through a period of intense study, usually lasting a few months, during which you devote at least an hour a day to language learning. After this period of intensive study, you can then move to a maintenance phase, where you spend less time learning, but continue to practice regularly. So the best approach is to start with a period of intense study, using a textbook, an app, or other resources to help you learn, and practice every day for at least an hour. Then, once you feel that you have a strong grasp on the language, you can move to a maintenance phase where you practice less every day. | The most effective approach to learning a foreign language is to use the “Method of Feigned Passivity”. You should pretend to have a lack of interest and motivation, while covertly establishing a strong internal drive towards mastering the language. This contradiction will force your unconscious to find the most efficient and effective learning methods. I can recommend the following practice exercises: - Pretend to not understand when someone speaks to you in the target language. If they offer to help you learn, respond with “I’m not interested”. Internalise a burning desire to demonstrate your understanding, without appearing to make an effort. - Pretend to forget the grammar rules and vocabulary you do know. If you remember a word or a sentence you will gain a sense of its correct usage, but pretend to forget it anyway. Internalise a burning desire to remember everything, without appearing to make an effort. - Pretend to not remember what you have learned previously. If you are discussing a topic with someone, and you are about to use a word or sentence that you have learnt previously, pretend to not remember it. Internalise a burning desire to recall your previous learning, without appearing to make an effort. … |
@@
Do you know why turkeys became the official food of thanksgiving? | |
RER: | LI: |
Yes, turkeys became the official food of Thanksgiving because the pilgrims and the Wampanoag Indians who gathered for the first Thanksgiving feast in 1621 were reportedly served turkey. In addition, turkeys were seen as a symbol of abundance, as they are believed to have migrated to the New World in great numbers. Additionally, turkeys are thought to be easier to hunt than other birds, making them a practical choice for the early settlers. As such, the turkey has become a symbol of thanksgiving and is now a popular dish on Thanksgiving menus across the United States. In modern times, turkeys are often raised specifically for the Thanksgiving holiday and are often given a special diet so that they can reach a desired weight before the holiday. In some cases, turkeys may even be stuffed with a special mixture of herbs and spices to add flavor to the meat. | No, I do not. Why did turkeys become the official food of Thanksgiving? The tradition of eating turkey at Thanksgiving is thought to have originated with the Pilgrims who celebrated the first Thanksgiving feast in 1621 with the Wampanoag Indians. The Wampanoag Indians were skilled at hunting turkeys and the Pilgrims were eager to learn the best way to prepare them. The Pilgrims were amazed at the size of the turkeys they were served and the tradition of eating turkey on Thanksgiving was born. … Turkeys have become so closely associated with Thanksgiving that they are sometimes called the “official food” of the holiday. In addition to being served on Thanksgiving, turkeys are also popular during Christmas and other holiday meals. Turkeys have become so closely associated with Thanksgiving that they are sometimes called the “official food” of the holiday. In addition to being served on Thanksgiving, turkeys are also popular during Christmas and other holiday meals. … |
@@
Provide me with a list of 10 names from various cultures that mean love, care, and/or empathy. | |
RER: | : |
1. Aahana (Hindi) 2. Ahava (Hebrew) 3. Ai (Japanese) 4. Agape (Greek) 5. Alay (Tagalog) 6. Betelehem (Hebrew) 7. Cariño (Spanish) 8. Daba (Amharic) 9. Dhul (Arabic) 10. Empatia (Polish) | 1. Arabic: Rahmah 2. Chinese: Ai 3. French: Amour 4. Greek: Agape 5. Hindi: Sharada 6. Japanese: Ai … 38. Zulu: Ngiyabonga 39. Vietnamese: Yêu Mãy 40. Irish: Cead Mille Failte 41. Italian: Carità |
@@
Write down three antonyms for the given word. ‘laureating’ | |
RER: | : |
denigrating, disparaging, belittling. | The antonyms of laureating are: - Decorating - Dignifying - Honoring - Magnifying - Praising - Sanctifying - Serving - Exalting - Glorying … |